OpenStudy (anonymous):

ok i have two problems but both have the same question. help please write the equation for the translation of y=4/x that has the given asymptotes. 1) x=0, y=3 2) x=-3, y= -4

OpenStudy (anonymous):

help @whpalmer4

OpenStudy (anonymous):

@jdoe0001

OpenStudy (whpalmer4):

We want to translate \(y = 4/x\) so that it has asymptotes \(x=0, y=3\).

OpenStudy (whpalmer4):

What are the asymptotes of \(y = 4/x\) before translation?

OpenStudy (anonymous):

its ok. i now this is smiler to the other problems but. i dont know how to work backwards.

OpenStudy (anonymous):

is it x=0 and y=0 ???? im still a little rusty on this asymptote stuff :)

OpenStudy (whpalmer4):

Yes. The fact that we have a 4 in the numerator instead of 1 doesn't change the asymptotes, because \(x\) is going to huge numbers (or very small numbers) and it washes out the effect of the 4. So if the one asymptote remains \(x=0\), are we shifting the graph up or down, or right or left?

OpenStudy (whpalmer4):

y = 3 is a horizontal or vertical line?

OpenStudy (anonymous):

sorry i said that wrong lol ummm what i was trying to say was is the graph going to move up to the 3 on the x axis?

OpenStudy (whpalmer4):

Yes, the graph will move up by 3 units. What do we have to do to \(y = f(x)\) to make it move up by 3 units?

OpenStudy (whpalmer4):

(doesn't matter what \(f(x)\) is, the process is the same)

OpenStudy (anonymous):

we will add 3

OpenStudy (anonymous):

right??

OpenStudy (whpalmer4):

Indeed we will! So what is our new function?

OpenStudy (anonymous):

is it y=4/x +3???

OpenStudy (whpalmer4):

Is that \[y = \frac{4}{x}+3\] or \[y = \frac{4+3}{x}\]?

OpenStudy (anonymous):

the first one.............is that right???? i hope so :/

OpenStudy (anonymous):

it will be the first one i graphed it to make sure. right??

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

yes, the first one is correct. Sorry, my sweetie is watching "Groundhog Day" and we just got to a particularly funny scene :-)

OpenStudy (anonymous):

oh ok its not a problem i understand.

OpenStudy (anonymous):

ok so for part 2) what will that one be. its harder than this first one

OpenStudy (anonymous):

it is x=-3 and y= -4

OpenStudy (anonymous):

i know we will subtract 4 from the equation. i know it will look like this if x was 0.......... y=4/x -4

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Okay, yes, subtracting 4 will shift the horizontal asymptote down to y = -4. Now we need to shift the graph sideways so the vertical asymptote is at x = -3. remember \(y=1/x\)? Where was the vertical asymptote? What happened there when you tried to compute \(y\)?

OpenStudy (anonymous):

aww man i dont remember.

OpenStudy (anonymous):

dang it!!!!! i thought i was doing so well

OpenStudy (whpalmer4):

The vertical asymptote was at \(x=0\), right? what happens when we plug \(x = 0\) into \(y = 1/x\)?

OpenStudy (anonymous):

yes ok now i remember. we didnt get anything there was a error

OpenStudy (whpalmer4):

Right. So we need to have an error like that at x = -3. How do we change our function to do that?

OpenStudy (anonymous):

im not sure how to do that

OpenStudy (whpalmer4):

what caused the error?

OpenStudy (anonymous):

when we divided by 0

OpenStudy (whpalmer4):

Okay, how do we divide by zero if x=-3?

OpenStudy (anonymous):

um i know if we that -3 and 3 cancel each other out??? if that helps any.....

OpenStudy (whpalmer4):

yes. it does. we want to have a 0 in the denominator when x = -3. \[x+a = 0\]\[x = -3\]\[-3+a = 0\] any ideas?

OpenStudy (anonymous):

a=3 right???

OpenStudy (whpalmer4):

Yes. so what is our new denominator? remember, it will equal 0 when x = -3

OpenStudy (anonymous):

umm -3+3

OpenStudy (whpalmer4):

can you write an expression involving \(x\) that =0 when x = -3?

OpenStudy (anonymous):

oh maybe 3x???

OpenStudy (whpalmer4):

3(-3) = -9, not 0

OpenStudy (whpalmer4):

just using addition or subtraction

OpenStudy (whpalmer4):

you're going to want to hit yourself in the forehead when you see the answer :-)

OpenStudy (anonymous):

dang i thought i had it then i tested it and saw i was wrong is it 0x????

OpenStudy (anonymous):

omg it will be x+3

OpenStudy (anonymous):

im a retard lol i did hit myself 0.0

OpenStudy (whpalmer4):

\[y=\frac{4}{x+3}-4\] What is the value of that at \(x = -3\)?

OpenStudy (anonymous):

-_- so bummed at myself

OpenStudy (whpalmer4):

"watch out forehead, here comes the palm!" :-)

OpenStudy (anonymous):

lol yep i should have known that

OpenStudy (whpalmer4):

so, to give the function a vertical asymptote at \(x = a\), just make sure there is a \((x-a)\) product term in the denominator

OpenStudy (anonymous):

wait what???

OpenStudy (whpalmer4):

we wanted one at \(x = -3\), so we needed \((x-(-3)) = (x+3)\) in the denominator.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

continue

OpenStudy (whpalmer4):

as a more general discussion, if we have a function \(y = f(x)\): \(y = f(x) + a\) shifts the graph up/down by \(a\) units \(y = af(x)\) expands the scale of the graph by a factor of \(a\) (and inverts it if \(a < 0\)) \(y = f(x+a)\) shifts the graph left/right by \(a\) units any or all of these can be combined to achieve multiple translations simultaneously

OpenStudy (anonymous):

oh ok i get it..... wel somewhat

OpenStudy (whpalmer4):

So we had \(y = 1/x\). Multiplying by \(a=4\) expanded the vertical scale by a factor of 4. Adding \(+3\) to the value of \(x\) (\(f(x+3))\) shifted the graph so that the vertical asymptote previously at \(x = 0\) is now at \(x = -3\).

OpenStudy (whpalmer4):

One more case I didn't mention: \(y = f(ax)\) expands/contracts the horizontal scale by a factor of \(a\)

OpenStudy (anonymous):

omg your going to kill me lol...... i have been doing this all day!!!!! my brain is fried lol 0.0 :S

OpenStudy (whpalmer4):

For the horizontal scale, if \(a < 1\) the graph spreads wider, if \(a > 1\) the graph is narrower. For vertical scale it works the other way. \(a>1\) makes the graph taller, \(a<1\) makes it shorter. Yeah, it's a bit overwhelming, isn't it :-)

OpenStudy (whpalmer4):

But save the page and come look at it again tomorrow and the next day.

OpenStudy (anonymous):

ok so is the answer y=4/x+3 -4 just want to be done until tomorrow. ill be back dont worry. lol =S

OpenStudy (whpalmer4):

Close: if you are going to write it horizontally like that, you MUST use parentheses to indicate that the denominator is (x+3). \[y = 4/(x+3) - 4\]

OpenStudy (anonymous):

ok i will thank you lol

OpenStudy (anonymous):

will you be on here tomorrow???

OpenStudy (whpalmer4):

Otherwise, the order of operator precedence says that we do multiplication and division before addition and subtraction, so \[y = 4/x+3 -4 \]turns into \[y = 4/x - 1 = -1 + 4/x\]and that isn't the same thing.

OpenStudy (whpalmer4):

Oh, I probably will be, though I couldn't tell you exactly when.

OpenStudy (anonymous):

ok well i will just keep a look out :D THANK YOU SO MUCH i couldnt have dont it without you =D

OpenStudy (whpalmer4):

If you send me a message, I'll get an email to that effect and know to come looking for you.

OpenStudy (anonymous):

ok i will do that. see you tomorrow